OpenStudy (anonymous):

Solve. Show all steps used to solve. 4(x - 6) = -2(x - 3) Solve. Show all steps used to solve. 4(x - 6) = -2(x - 3) @Mathematics

OpenStudy (anonymous):

distributiive property 4x-24=-2x-6 combine like terms 6x-24=6 6x=6+24 x=30/2 x=15